# JavisDiT [Uncaptioned image]: Joint Audio-Video Diffusion Transformer with Hierarchical Spatio-Temporal Prior Synchronization

### v1

@@ Abstract
This paper introduces JavisDiT, a novel Joint Audio-Video Diffusion Transformer designed for synchronized audio-video generation (JAVG). Built upon the powerful Diffusion Transformer (DiT) architecture, JavisDiT is able to generate high-quality audio and video content simultaneously from open-ended user prompts. To ensure optimal synchronization, we introduce a fine-grained spatio-temporal alignment mechanism through a Hierarchical Spatial-Temporal Synchronized Prior (HiST-Sypo) Estimator. This module extracts both global and fine-grained spatio-temporal priors, guiding the synchronization between the visual and auditory components. Furthermore, we propose a new benchmark, JavisBench, consisting of 10,140 high-quality text-captioned sounding videos spanning diverse scenes and complex real-world scenarios. Further, we specifically devise a robust metric for evaluating the synchronization between generated audio-video pairs in real-world complex content. Experimental results demonstrate that JavisDiT significantly outperforms existing methods by ensuring both high-quality generation and precise synchronization, setting a new standard for JAVG tasks. Our code, model, and dataset will be made publicly available at https://javisdit.github.io/.
@@ 1 Introduction
Figure 1: Given the input text prompt, a JAVG system generates a spatial-temporally synchronized sounding video. The sounds align perfectly with the temporal progression of the actions.
@@
In the field of AI-generated content (AIGC), the generation of multimodal content, such as images, videos, and audio, has recently garnered unprecedented attention [41, 27, 59, 36], where diffusion-based models [13, 29, 34] have demonstrated remarkable performance. While earlier works primarily focused on generating single-modal content, there is now a growing interest in simultaneously generating multiple modalities [9, 54]. Notably, synchronized audio and video generation [42, 56, 48] has emerged as a crucial area of study. Audio and video are inherently interconnected in most real-world scenarios, making their joint generation highly valuable for applications, such as movie production and short video creation. Current approaches to synchronized audio-video generation can be broadly categorized into two types. The first involves asynchronous pipelines, where audio is generated first and then used to synthesize video [60, 18, 62], or vice versa [6, 63, 55]. The second type involves end-to-end Joint Audio-Video Generation (namely JAVG) [42, 56, 48], which has attracted more research attention by avoiding cascaded noise accumulation. Overall, it is strongly and widely believed that achieving high-quality JAVG requires two equally critical criteria: (1) ensuring high-quality generation of audio and video, and (2) maintaining perfect synchronization between two modalities.
Figure 2: On the left, we present the overall DiT-based sounding video generation architecture of our JavisDiT Refer to caption System, consisting of a video generation branch, audio generation branch, and the MM-BiCrossAttn module. On the right, we illustrate the detailed structural design of Spatio-temporal Self-attention (ST Self-Attn), Fine-grained Spatio-temporal Cross-attention (Fine-grained ST-CrossAttn), and Multi-Modality Bidirectional Cross-attention (MM-BiCrossAttn).
@@
On one hand, for high-quality content generation of audio and video, the backbone architectures of both modules need to be as strong as possible. Benefiting from the superior generative capabilities of diffusion models, most current JAVG research [42, 56, 48] adopts diffusion backbone architectures. Recently, the Diffusion Transformer (DiT) [34] architecture has demonstrated remarkably enhanced performance for vision and even audio tasks, also offering advantages in efficiency and scalability for concurrent works [51, 48, 26, 64]. In particular, AV-DiT [51] and MM-LDM [48] inherit an image DiT [34, 8] to generate video and audios, where the fine-grained spatio-temporal modeling capability can be however limited. Uniform [64] and SyncFlow [26] take the enhanced STDiT3 [65] blocks, but they either simply concatenate video/audio tokens or utilize a mono-directional video-to-audio adapter, lacking sufficient mutual information exchange between two channels.
@@
On the other hand, in terms of synchronization between the audio and video channels in JAVG, extensive investigations have been made to explore more effective alignment strategies. However, most of these works still fall prey to insufficient modeling of synchronization. For example, some emphasize coarse-grained temporal alignment by implementing straightforward parameter sharing [42, 51, 64] or temporal modulating [51, 26] between the audio and video channels. Others focus on coarse-grained semantic alignment, such as Seeing-Hearing [56], which carries out simple audio-video embedding alignment, and MM-LDM [48] performing straightforward representation alignment in the latent space of the diffusion model. These approaches fail to consider a fine-grained spatio-temporal alignment, which is essential for realistic synchronized audio-video generation. Specifically, in a realistic-sounding video, all visual and auditory contents should be dictated by spatio-temporal characteristics of various objects, i.e., a) spatially distinct visual content (e.g., where events occur and how objects move), and b) temporally corresponding auditory attributes, such as timbre and synchronized duration. We illustrate this in Fig. 1: A dog and a robot are playing on the ground, while an alien appears and starts to talk. In this scene, the robot’s mechanical whirs and the dog’s squeaks and barks keep consistent, and the alien’s talking emerges later and continues to the end. Overall, the visual content and audio stay synchronized both temporally and spatially.
@@
To address the above dual challenges, this paper proposes a novel niche-targeting JAVG system, called JavisDiT (cf. Fig. 2). First, JavisDiT adopts the DiT architecture as the backbone, where the audio and video channels share AV-DiT blocks to enable high-quality audio-video generation. Within JavisDiT, we design three infrastructural blocks: Spatio-Temporal Self-Attention, Coarse-Grained Cross-Attention, and Fine-Grained Spatio-Temporal Self-Attention Cross-Attention. To implement the fine-grained spatio-temporal alignment idea, we design a Hierarchical Spatial-Temporal Synchronized Prior (HiST-Sypo) estimation module. The HiST-Sypo Estimator hierarchically extracts the following from the input conditional prompt:
$\bullet$ global coarse-grained spatio-temporal priors, such as the semantic framework of the overall sounding video;
$\bullet$ fine-grained spatio-temporal priors, such as the distinct visual content triggered by various sounds and their corresponding temporal priors.
@@
These global and fine-grained priors serve as spatio-temporal features, injected into different AV-DiT blocks to guide the spatial semantic and temporal synchronization of both audio and video. Building on JavisDiT, we further design a contrastive learning-based [3] HiST-Sypo estimation strategy to learn robust spatial-temporal synchronized prior knowledge from large-scale data of sounding videos.
@@
Further, we observe that the existing JAVG benchmarks, e.g., Landscape [21] and AIST++ [22] suffer from certain limitations, including overly simplistic audio-video content and limited scene diversity [31]. This creates an evident gap compared to the complex audio-video content encountered in real-world applications, i.e., models [42, 48] trained there largely suffer from out-of-domain issue, thereby partially hindering the development and practical applicability on open environment. To bridge these gaps, we introduce a new JAVG benchmark, called JavisBench. After rigorous manual inspection, we collect a total of 10,140 high-quality text-captioned sounding videos, which span over 5 dimensions with 19 scene categories. Notably, more than 50% of these videos feature highly complex and challenging scenarios, making the dataset more representative of diverse real-world applications. Meanwhile, we find that existing JAVG evaluation method can also be limited by failing to adequately assess the JAVG systems on complex sounding videos. Thus, we devise a novel metric (JavisScore) based on a temporal-aware semantic alignment mechanism.
@@
  • We introduce JavisDiT, a novel DiT-based JAVG system with a hierarchical spatio-temporal prior estimation mechanism to achieve synchronized audio-video alignment in both spatial and temporal dimensions.
@@ 2 Preliminaries
Task Definition and Formulation. Joint Audio-Video Generation (JAVG) requires diffusion models $\mathcal{G}_{\theta}$ to simultaneously generate videos $\mathbf{v}$ and corresponding audios $\mathbf{a}$ given a text input $s$. The forward process at timestamp $t$, either with DDPM [13] or FlowMatching [24], is formulated as:
@@
In general practice, $\mathbf{v}\in\mathbb{R}^{T_{v}\times(H\times W)\times 3}$ denotes a video with $T_{v}$ frames and 3 RGB channels, where $H,W$ are the height and width of each video frame. On the other hand, $\mathbf{a}\in\mathbb{R}^{T_{a}\times M\times 1}$ encodes audios with the image-like MelSpectrogram with $T_{a}$ temporal frames and 1 channel, where $M$ is the frequency bins of the audio spectrogram.
@@
This work focuses on the joint generation of high-quality and spatio-temporally aligned video-audio pairs.
3 The Proposed JavisDiT [Uncaptioned image] System
@@ 3.1 JavisDiT Model Architecture
Spatio-Temporal Self-Attention. Given that both videos and audios possess spatial and temporal attributes [42], we employ a cascaded spatio-temporal self-attention mechanism for intra-modal information aggregation. As Fig. 2(b) illustrates, MHSA is applied sequentially along the spatial ($(H\times W)$ for $\mathbf{v}$, $M$ for $\mathbf{a}$) and temporal ($T_{v}$ for $\mathbf{v}$, $T_{a}$ for $\mathbf{a}$) dimensions. This efficiently achieves fine-grained spatio-temporal modeling with reduced computational cost.
@@
Spatio-Temporal Cross-Attention. For a text prompt $s$, we estimate spatial and temporal priors with $N_{s}$ and $N_{t}$ learnable $c$-dimensional tokens by our ST-Prior Estimator (details are presented in Sec. 3.2). As shown in Fig. 2(b), spatial and temporal priors guide cross-attention in both branches along the spatial and temporal dimensions, enabling unified, fine-grained conditioning for video-audio synchronization.
@@
Cross-Modality Bidirectional-Attention. After aligning video and audio with ST-Prior, we incorporate a bidirectional attention module [28] to enable direct cross-modal interactions. As Fig. 2(b) illustrates, after computing the attention matrix $A$ between $q_{v}$ and $k_{a}$, we first multiply $A$ with $v_{a}$ to obtain the audio-to-video cross-attention. Similarly, multiplying $A^{T}$ (the transpose of $A$) with $v_{v}$ yields the video-to-audio cross-attention. This mechanism enhances cross-modal information aggregation to facilitate high-quality joint audio-video generation.
3.2 Hierarchical Spatio-Temporal Prior
@@
Unlike previous works, our JavisDiT derives two hierarchical conditions (priors) from texts: a global semantic prior for coarse event suggestion (what) and a fine-grained spatio-temporal prior (ST-Prior) to specify event timing and location (when and where), enabling precise synchronization between generated video and audio.
@@
Coarse-Grained Spatio-Temporal Prior. Since the default semantic embeddings from T5 encoder [37] are strong enough to coarsely describe the overall sounding event, we simply reuse T5 embeddings as our coarse-grained spatio-temporal prior (or, semantic prior).
@@
Fine-Grained Spatio-Temporal Prior. Text inputs typically provide coarse-grained descriptions of events, e.g., “a car starts its engine and leaves the screen”. To enable fine-grained conditioning, we estimate spatio-temporal priors: spatial prior specifies where events occur (e.g., “the car is in the top-left corner of screen”), and temporal prior defines when they start and end (e.g., “sound starts at 2s, exits at 7s, fades at 9s”). Instead of generating explicit prompts from LLMs [27], we efficiently estimate spatio-temporal priors as latent token conditions to guide the diffusion process.
@@
Figure 3: The framework of Spatio-temporal Prior Estimator with a 4-layer transformer encoder-decoder (referring to the purple region). Contrastive learning is utilized to optimize the estimator.
@@
Fine-Grained ST-Prior Estimation. As shown in Fig. 3, for a given text input $s$, we utilize the 77 hidden states from ImageBind’s [11] text encoder, and use $N_{s}=32$ spatial tokens $\mathbf{p}_{s}$ and $N_{t}=32$ temporal tokens $\mathbf{p}_{t}$ to query a 4-layer transformer $\mathcal{P}$ to extract spatio-temporal information. Since input text prompts usually may not specify an event’s happening time and location, it may almost occur and cease in arbitrary location and timing, and the same text $s$ should correspondingly yield different ST-Priors $(\mathbf{p}_{s},\mathbf{p}_{t})$. To capture this variability, our ST-Prior Estimator $\mathcal{P}$ outputs the mean and variance of a Gaussian distribution, from which a plausible $(\mathbf{p}_{s},\mathbf{p}_{t})$ is sampled: $(\mathbf{p}_{s},\mathbf{p}_{t})\leftarrow\mathcal{P}_{\phi}(s;\epsilon)$, where $\epsilon$ is a Gaussian noise. Furthermore, we carefully devised a contrastive learning approach to learn a robust ST-Prior Estimator, which involves a series of negative sample (asynchronous video-audio pairs) construction strategies and specifically-designed loss functions. Implementation details are provided in the Sec. C.2.

### v2

# Fan Zhou1,  Jiebo Luo4,  Ziwei Liu5,  Hao Fei2,  Tat-Seng Chua2

@@ Abstract
This paper introduces JavisDiT, a novel Joint Audio-Video Diffusion Transformer designed for synchronized audio-video generation (JAVG). Based on the powerful Diffusion Transformer (DiT) architecture, JavisDiT simultaneously generates high-quality audio and video content from open-ended user prompts in a unified framework. To ensure audio-video synchronization, we introduce a fine-grained spatio-temporal alignment mechanism through a Hierarchical Spatial-Temporal Synchronized Prior (HiST-Sypo) Estimator. This module extracts both global and fine-grained spatio-temporal priors, guiding the synchronization between the visual and auditory components. Furthermore, we propose a new benchmark, JavisBench, which consists of 10,140 high-quality text-captioned sounding videos and focuses on synchronization evaluation in diverse and complex real-world scenarios. Further, we specifically devise a robust metric for measuring the synchrony between generated audio-video pairs in real-world content. Experimental results demonstrate that JavisDiT significantly outperforms existing methods by ensuring both high-quality generation and precise synchronization, setting a new standard for JAVG tasks. Our code, model, and data are available at https://javisverse.github.io/JavisDiT-page/.
@@ 1 Introduction
In the field of AI-generated content (AIGC), multimodal generation that covers images, videos, and audio has gained increasing attention (Lyu et al., 2024; Liu et al., 2024b; Yang et al., 2024b; Lyu et al., 2025), where diffusion-based models (Ho et al., 2020; Liu et al., 2023b; Peebles & Xie, 2023) have demonstrated remarkable performance. While early studies mainly focused on single-modality generation, recent work has shifted toward the simultaneous generation of multiple modalities (Gadre et al., 2024; Wu et al., ). Notably, synchronized audio and video generation (Ruan et al., 2023; Xing et al., 2024; Sun et al., 2024) has emerged as a crucial area of study. Audio and video are inherently interconnected in most real-world scenarios, making their joint generation highly valuable for applications, such as movie production and short video creation. Current approaches to synchronized audio-video generation can be broadly categorized into two types. The first involves asynchronous pipelines, where audio is generated first and then used to synthesize video (Yariv et al., 2024; Jeong et al., 2023; Zhang et al., 2025), or vice versa (Comunità et al., 2024; Zhang et al., 2024; Xie et al., 2024). The second type involves end-to-end Joint Audio-Video Generation (namely JAVG) (Ruan et al., 2023; Xing et al., 2024; Sun et al., 2024), which, by avoiding noise accumulation during cascading processes of the first type, has thus attracted more research attention. Overall, it is strongly and widely believed that promoting JAVG requires two equally critical criteria: (1) ensuring high-quality generation of audio and video, and (2) pursuing precise synchronization between two modalities.
@@
On one hand, for high-quality audio-video generation, the backbone modules of both branches need to be as strong as possible. Recent JAVG research (Ruan et al., 2023; Xing et al., 2024; Sun et al., 2024) attempts to take Diffusion Transformer (DiT) (Peebles & Xie, 2023) as backbone architectures, due to the remarkably enhanced performance for vision and even audio tasks. In particular, AV-DiT (Wang et al., 2024b) and MM-LDM (Sun et al., 2024) inherit an image DiT (Peebles & Xie, 2023; Esser et al., 2024) to generate video and audio, where the fine-grained spatio-temporal modeling capability can be however limited. Concurrent works like Uniform (Zhao et al., 2025) and SyncFlow (Liu et al., 2024a) take the enhanced STDiT3 (Zheng et al., 2024) blocks, but they either simply concatenate video/audio tokens or utilize a mono-directional video-to-audio adapter, lacking sufficient mutual information exchange between the two channels.
Figure 1: Given the text prompt, a JAVG system generates a spatial-temporally synchronized sounding video. The sounds align precisely with the temporal progression of the actions.
@@
On the other hand, in terms of synchronization between the audio and video channels in JAVG, extensive investigations have been made to explore more effective alignment strategies. However, most of these works still fall prey to insufficient modeling of synchronization. For example, some emphasize coarse-grained temporal alignment by implementing straightforward parameter sharing (Ruan et al., 2023; Wang et al., 2024b; Zhao et al., 2025) or temporal modulating (Wang et al., 2024b; Liu et al., 2024a) between the audio and video channels. Others focus on coarse-grained semantic alignment, such as Seeing-Hearing (Xing et al., 2024), which carries out simple audio-video embedding alignment, and MM-LDM (Sun et al., 2024) performing straightforward representation alignment in the latent space of the diffusion model. These approaches fail to consider a fine-grained spatio-temporal alignment, which is essential for realistic synchronized audio-video generation. Specifically, in a realistic-sounding video, all visual and auditory contents should be dictated by spatio-temporal characteristics of various objects, i.e., a) spatially distinct visual content (e.g., where events occur and how objects move), and b) temporally corresponding auditory attributes, such as source category and synchronized duration. We illustrate this in Fig. 1: A dog and a robot are playing on the ground, while an alien appears and starts to talk. In this scene, the robot’s mechanical whirs and the dog’s squeaks and barks keep consistent, and the alien’s talking emerges later and continues to the end. Overall, the visual content and audio stay synchronized both temporally and spatially.
@@
To address the above dual challenges, this paper proposes a novel niche-targeting JAVG system, called JavisDiT (cf. Fig. 2). First, JavisDiT adopts the DiT architecture as the backbone, where the audio and video channels share AV-DiT blocks to enable high-quality audio-video generation. Within JavisDiT, we design three infrastructural blocks: Spatio-Temporal Self-Attention, Coarse-Grained Cross-Attention, and Fine-Grained Spatio-Temporal Cross-Attention (ST-CorssAttn). To implement the fine-grained spatio-temporal alignment idea, we design a Hierarchical Spatial-Temporal Synchronized Prior (HiST-Sypo) estimation module. The HiST-Sypo Estimator hierarchically extracts the following from the input conditional prompt: $\bullet$ global coarse-grained spatio-temporal priors, such as the semantic framework of the overall sounding video; $\bullet$ fine-grained spatio-temporal priors, such as the distinct visual content triggered by various sounds and their corresponding temporal priors.
@@
These global and fine-grained priors serve as spatio-temporal features, injected into different AV-DiT blocks to guide the spatial semantic and temporal synchronization of both audio and video. Built on JavisDiT, we further design a contrastive learning-based (Chen et al., 2020b) HiST-Sypo estimation strategy to learn robust spatial-temporal synchronized prior knowledge from large-scale data of sounding videos.
@@
Further, we observe that the existing JAVG evaluation benchmarks, e.g., Landscape (Lee et al., 2022) and AIST++ (Li et al., 2021) suffer from certain limitations, including overly simplistic audio-video content and limited scene diversity (Mao et al., 2024). This creates an evident gap compared to the complex audio-video content encountered in real-world applications, i.e., models (Ruan et al., 2023; Sun et al., 2024) trained there largely suffer from out-of-domain issue, thereby partially hindering the development and practical applicability on open environment. To bridge these gaps, we introduce a new JAVG benchmark, called JavisBench. After rigorous manual inspection, we collect a total of 10,140 high-quality text-captioned sounding videos, which span over 5 dimensions with 19 scene categories. Notably, more than 50% of these videos feature highly complex and challenging scenarios, making the dataset more representative of diverse real-world applications. Meanwhile, we find that existing JAVG evaluation method can also be limited by failing to adequately assess the JAVG systems on complex sounding videos. Thus, we devise a novel metric (JavisScore) based on a temporal-aware semantic alignment mechanism.
@@
  • We introduce JavisDiT, a novel JAVG model with a hierarchical spatio-temporal prior estimation mechanism to achieve audio-video synchronization in both spatial and temporal dimensions.
@@ 2 Preliminaries
Task Definition and Formulation. Joint Audio-Video Generation (JAVG) requires diffusion models $\mathcal{G}_{\theta}$ to simultaneously generate videos $\mathbf{v}$ and corresponding audios $\mathbf{a}$ given a text input $s$. The forward process at timestamp $t$, either with DDPM (Ho et al., 2020) or FlowMatching (Lipman et al., 2023), is formulated as:
@@
$\mathbf{v}\in\mathbb{R}^{T_{v}\times(H\times W)\times 3}$ denotes a video with $T_{v}$ frames and 3 RGB channels, where $H,W$ are the frame height and width. On the other hand, $\mathbf{a}\in\mathbb{R}^{T_{a}\times M\times 1}$ encodes audio in the image-like mel-spectrogram with $T_{a}$ temporal frames and 1 channel, where $M$ is the frequency bins.
@@
We focus on the joint generation of high-quality and spatio-temporally aligned video-audio pairs.
Figure 2: On the left, we present the overall DiT-based sounding video generation architecture of JavisDiT, consisting of a video generation branch, audio generation branch, HiST-Sypo Estimator module, and the $M$-BiCrossAttn module. On the right, we illustrate the detailed structural design of Spatio-temporal Self-attention (ST-SelfAttn), Fine-grained Spatio-temporal Cross-attention (Fine-Grained ST-CrossAttn), and Multi-Modality Bidirectional Cross-attention ($M$-BiCrossAttn).
3 The Proposed JavisDiT System
@@ 3.1 JavisDiT Model Architecture
Spatio-Temporal Self-Attention. Given that both videos and audios possess spatial and temporal attributes (Ruan et al., 2023), we employ a cascaded spatio-temporal self-attention mechanism for intra-modal information aggregation. As illustrated Fig. 2(b), MHSA is applied sequentially along the spatial ($(H\times W)$ for $\mathbf{v}$, $M$ for $\mathbf{a}$) and temporal ($T_{v}$ for $\mathbf{v}$, $T_{a}$ for $\mathbf{a}$) dimensions. This efficiently achieves fine-grained spatio-temporal modeling with reduced computational cost.
@@
Spatio-Temporal Cross-Attention. For a text prompt $s$, after injecting coarse semantics from the T5 encoder (Raffel et al., 2020) via vanilla cross attention, we estimate spatial and temporal priors with $N_{s}$ and $N_{t}$ learnable $c$-dimensional tokens by our ST-Prior Estimator (see details in Sec. 3.2). As Fig. 2(b) shows, spatial and temporal priors guide cross-attention in both branches along the spatial and temporal dimensions, enabling unified, fine-grained conditioning for video-audio synchronization.
@@
Cross-Modality Bidirectional-Attention. After aligning video and audio with ST-Prior, we incorporate a bidirectional attention module (Liu et al., 2025) to enable direct cross-modal interactions. As Fig. 2(b) illustrates, after computing the attention matrix $A$ between $q_{v}$ and $k_{a}$, we first multiply $A$ with $v_{a}$ to obtain the audio-to-video cross-attention. Similarly, multiplying $A^{T}$ (the transpose of $A$) with $v_{v}$ yields the video-to-audio cross-attention. This mechanism enhances cross-modal information aggregation to facilitate high-quality joint audio-video generation.
3.2 Hierarchical Spatial-Temporal Synchronized Prior Estimator
@@
Unlike previous works, our JavisDiT derives two hierarchical conditions (priors) from texts: a global semantic prior for coarse event suggestion (what) and a fine-grained spatio-temporal prior (ST-Prior) to specify event timing and location (when and where). These two complementary priors are unified into a HiST-Sypo Estimator, enabling precise synchronization between generated video and audio.
@@
Coarse-Grained Spatio-Temporal Prior. Since the default semantic embeddings from T5 encoder (Raffel et al., 2020) are strong enough to coarsely describe the overall sounding event, we simply reuse T5 embeddings as our coarse-grained spatio-temporal prior (or, semantic prior).
@@
Fine-Grained Spatio-Temporal Prior. Text inputs typically provide coarse-grained descriptions of events, e.g., “a car starts its engine and leaves the screen”. To enable fine-grained conditioning, we estimate spatio-temporal priors: spatial prior specifies where events occur (e.g., “the car is in the top-left corner of screen”), and temporal prior defines when they start and end (e.g., “sound starts at 2s, exits at 7s, fades at 9s”). Instead of generating explicit prompts from LLMs (Liu et al., 2024b), we efficiently estimate spatio-temporal priors as latent token conditions to guide the diffusion process.
@@
Figure 3: The framework of Spatio-temporal Prior Estimator with a 4-layer transformer encoder-decoder (referring to the purple region). Contrastive learning is utilized for optimization.
@@
Fine-Grained ST-Prior Estimation. As shown in Fig. 3, for a given text input $s$, we utilize the 77 hidden states from ImageBind’s (Girdhar et al., 2023) text encoder, and use $N_{s}=32$ spatial tokens $\mathbf{p}_{s}$ and $N_{t}=32$ temporal tokens $\mathbf{p}_{t}$ to query a 4-layer transformer $\mathcal{P}$ to extract spatio-temporal information. Since input text prompts usually may not specify an event’s happening time and location, it may almost occur and cease in arbitrary location and timing, and the same text $s$ should correspondingly yield different ST-Priors $(\mathbf{p}_{s},\mathbf{p}_{t})$. To capture this variability, our ST-Prior Estimator $\mathcal{P}$ outputs the mean and variance of a Gaussian distribution, from which a plausible $(\mathbf{p}_{s},\mathbf{p}_{t})$ is sampled: $(\mathbf{p}_{s},\mathbf{p}_{t})\leftarrow\mathcal{P}_{\phi}(s;\epsilon)$, where $\epsilon$ is a Gaussian noise. Furthermore, we carefully devised a contrastive learning approach to learn a robust ST-Prior Estimator, which involves a series of negative sample (asynchronous video-audio pairs) construction strategies and specifically-designed loss functions. Implementation details are provided in the Sec. C.2.